OpenStudy (anonymous):

Hey guys can you help me solve equation x-1/9=2? Thanks

OpenStudy (anonymous):

x=2 and 1/9

OpenStudy (anonymous):

yeah that what i was thinking but it shows that you can choose from 18, 17, 12 and 19

OpenStudy (anonymous):

no problem all you had to do was add 1/9 to the other side of the equation and cancel it out on the side that it was on :)

OpenStudy (anonymous):

so what is the answer? :)

OpenStudy (phi):

is this \[ \frac{x-1}{9}=2 \]?

OpenStudy (anonymous):

yeah :)phi

OpenStudy (phi):

multiply both sides by 9 what do you get?

OpenStudy (phi):

\[ 9 \cdot \frac{x-1}{9}= 2 \cdot 9\]

OpenStudy (anonymous):

so the answer is 17

OpenStudy (anonymous):

Thank You both :)

OpenStudy (anonymous):

answer is 19!! as phil showed if x-1 = 18 then x = 19 not 17

OpenStudy (phi):

you can always try to check the answer by subbing in for x in the original equation (x-1)/9 with x=19 gives us (19-1)/9 or 18/9 or 2